a protein smaple (mw2500) having concentration of 0.5ug/ml and extinction coefficient of 25000dm cubic /mol/cm and absorbs uv light. calculate the percentage of light transmitted through this sample if path length is 1cm
1
Expert's answer
2020-04-06T11:58:58-0400

According to Lambert-Beer's law:

%T = 100% × 10-A

A = ε × c × l

с = n/V = m/(Mr × V)

where %T - percentage transmittance, A - absorbance, ε - molar extinction coefficient, c - concentration, l - path length, Mr - molecular weight, m - weight, V - volume, n - number of moles.

The final equation is as follwoing:

%T = 100% × 10-ε × l × m/(Mr × V)

%T(protein) = 100% ×10-25000 × 1 cm × 0.5 ug / (2500 g/mol × 1 ml) = 99%


Answer: 99%